OpenStudy (anonymous):

The annual yield per lemon tree is fairly constant at 320 pounds when the number of trees per acre is 50 or fewer. for each additional tree over 50, the annual yield per tree for all trees on the acre decreases by 4 pounds due to overcrowding. find the number of trees that should be planted on an acre to produced the maximum yield. how many pounds is the maximum yield?

OpenStudy (anonymous):

have i become lik ur personal answerer? :P jkjk

OpenStudy (anonymous):

lmfao!! naw is because i trust your answers & the other people nope>.<

OpenStudy (anonymous):

lololol ... thanks XD

OpenStudy (anonymous):

you're welcome^-^

OpenStudy (anonymous):

Let x = no. of trees, over 50, planted Then y = total yield per acre : total yield = No.of trees * lbs of lemons per tree : y = (50+x) * (320 - 4x) y = 16000 - 200x + 320x - 4x^2 The quadratic equation y = -4x^2 + 120x + 16000 : The axis of symmetry will be the value of x for max y The formula for that: x = -b/2a In this problem: a=-4; b=120

OpenStudy (anonymous):

is this kinda right?

OpenStudy (anonymous):

yeahhh thats what i was going for... somehow i got caught up in something tho... XD.... prob cuz its getting late XD

OpenStudy (anonymous):

true>.< but i need to finish this so i could do my physics paper next>.< but this is my last problem so thats good but help me dude im stuck>.<

OpenStudy (anonymous):

ok so u get ur x from \[x = \frac{-b}{2a}\] to be 15 right?

OpenStudy (anonymous):

so then once u get that... just plug it into the equation of \[y=-4x^{2} +120x +16000\]

OpenStudy (anonymous):

and thats it?

OpenStudy (anonymous):

i would think... hold on... lemme check..

OpenStudy (anonymous):

yup thats ur max

OpenStudy (anonymous):

ya can get me bak if im wrong

OpenStudy (anonymous):

wait so the #'s of trees planted would be 15 right and the max is 16900?

OpenStudy (anonymous):

well the number of trees planted would be 15+ 50 gibing you 65

OpenStudy (anonymous):

and the max is correct XD

OpenStudy (anonymous):

you positive>.<

OpenStudy (anonymous):

if 15 is the definite most correct answer you can get... then yes... :D

OpenStudy (anonymous):

lol dude im confused in the find the number of trees that should be planted on an acre to produced the maximum yield part >.<